Calculate the pH after the addition of each of the following volumes when 20.00 mL of 0.125 M NaOCl is titrated with 0.250M HCl. (Ka = 3.0 * 10^-8 for HOCl) (Hint: This is a weak base-strong acid titration)

A) 5.0 mL of HCl added

B)10.0 mL of HCl added

C) 17.0 mL of HCl added
Answer #1

Given that concentration of HCl = 0.250 M concentration of NaClO = 0.125 M Volume of NaClo= 20.00 mL Ka for HClO = 3.0 × 10-8

a) Addition of 5.00 mL of HCI. Number of moles of NaClO = M*V = 0.125 M* 20.00 mL = 2.5 mmol Number of moles of HCl = M*V = 0

(b) Addition of 10.00 mL of HCI Number of moles of NaClO = M*V = 0.125 M* 20.00 mL = 2.5 mmol Number of moles of HCI = M*V =

x.x Kar -= 3.00 x10-8 0.083333-X x is small, so 0.083333-X=0.083333 X.X 0.083333 =3.00x10-8 3x2 = 2.50 x10-9 x=5.00x10--M ::[